Write an equation of the cosine function with amplitude 2 and period 8π.
1
Expert's answer
2012-02-07T13:02:43-0500
f(x)=2cos(x/4)
Let's verify if written function meets the requirements:
|2cos(x/4)|<=2& -amplitude
f(x+8pi)=2cos((x+8pi)/4)=2cos(x/4+2pi)=2cos(x/4)=f(x) - so 8pi is the period